OpenStudy (anonymous):

Simplify and state the excluded values. 24n^7 / 24n^4 Any help would be greatly appreciated. :)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{24n^7}{24n^4} }\) well the first thing here would be to CANCEL 24 on top and bottom. \(\huge\color{blue}{ \frac{\cancel{ 24 } n^7}{\cancel{ 24 }n^4} }\) \(\huge\color{blue}{ \frac{n^7}{n^4} }\) (solving this more ...) \(\huge\color{blue}{ n^7 \div n^4=n^{(7-4)}=n^3 }\) (b/c you subtract exponents when dividing such numbers) MAKES SENSE ?

OpenStudy (anonymous):

So would it be: